A Bibliography of Publications in the ALGOL Bulletin

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 May 2026 Version 1.01

Title word cross-reference

n [445].

-cycle [445].

10 [319]. 10.1.3 [28]. 1959 [8, 20]. 1960 [75]. 1961 [75].

2 [352]. 2.1 [344]. 200 [391]. 205 [474]. 2900 [434].

3 [359].

4 [369].

60 [71, 64, 175, 178, 193, 54, 16, 53, 56, 145, 134, 287, 360, 386, 456, 19, 27, 39, 61, 291, 15, 365, 111, 194, 308]. 68’ [442, 265, 328, 384, 398, 339, 315, 314, 289, 436, 469, 366, 435, 358, 320, 427, 474, 404, 375, 376, 395, 394, 326, 348, 361, 266, 452, 385, 429, 430, 418, 340, 357, 447, 292, 325, 338, 344, 367, 397, 409, 437, 425, 424, 434, 443, 444, 457, 468, 467, 473, 334, 330, 301, 310, 464, 288, 351, 400, 419, 428, 356, 349, 374, 458, 446, 475, 381, 329, 273, 416, 234, 438]. 68C [459, 425]. 6x [148].

9 [318].

A. [149, 137]. AB14 [90]. AB19.3.10 [157]. AB19.3.11.4 [149]. AB21.3.6 [201, 210]. AB24.3.4 [229]. abbreviations [231, 243]. ABSTRACTd [412]. Abstracto [401, 402]. ACM [151, 122]. active [309]. Activity [304]. Actual [84, 77]. Actual-Formal [84]. Adaption [328]. Additional [209]. Addressing [179, 177]. ALCOR [20, 36]. Algebraic [429]. ALGOL [162, 396, 208, 212, 200, 442, 75, 464, 230, 265, 328, 71, 162, 221, 384, 398, 339, 350, 315, 314, 289, 110, 119, 436, 469, 366, 64, 355, 103, 175, 435, 459, 358, 320, 185, 243, 427, 474, 404, 375, 376, 395, 394, 326, 348, 361, 178, 193, 266, 76, 452, 385, 278, 429, 430, 56, 418, 145, 134, 287, 192, 340, 357, 447, 292, 325, 338, 360, 352, 359, 344, 367, 369, 386, 397, 409, 437, 425, 424, 434, 444, 456, 457, 468, 467, 473, 128, 334, 186, 330, 8, 14, 11, 13, 19, 27, 39, 17, 35, 51]. ALGOL [52, 22, 23, 26, 34, 25, 37, 68, 73, 81, 59, 69, 70, 88, 87, 89, 96, 97, 98, 102, 253, 301, 310, 291, 288, 351, 400, 124, 419, 277, 428, 40, 148, 171, 356, 341, 349, 374, 458, 446, 475, 381, 383, 329, 240, 365, 111, 209, 263, 237, 194, 416, 58, 234, 308, 172, 471, 54, 16, 53, 61, 15, 273, 438, 443]. Algol-60 [16]. ALGOL-like [69]. ALGOL60 [152, 118, 104, 163, 141, 165, 147]. ALGOL68 [256, 251, 252, 280]. Algollers [380]. Algorithm [402, 430]. Algorithmic [234, 438, 265]. algorithms [21]. Allocation [56, 15]. Allowing [190]. ALPHA [144]. amendment [252]. Analysis [400, 58]. Analyzer [418]. animals [150]. Announcements [324, 336, 343, 354, 363, 371, 378, 389, 406, 414, 422, 432, 440, 449, 454, 461, 466, 472, 100, 107, 115, 156, 169, 183, 198, 205, 225, 216, 236, 249, 285, 306]. applied [243]. Arrays [15, 142, 67]. Art [462]. ASERC [256]. aspects [321]. Assignment [164, 121]. Association [60]. Atoms [290]. August [75]. Automata [331]. Automatic [55, 144].

B [450]. back [38, 24]. Balancing [339]. Basic [221, 256]. Beginner [410]. Between [303]. Blemishes [141]. Blocks [16, 79]. bodies [50, 78, 79]. boy [109, 136]. brief [349]. Browse [476]. Bulletin [13, 52, 22, 73, 81, 87, 96, 471, 14, 89]. Bulletins [476].

CACM [21]. called [95, 41]. Case [123, 132]. CDC [464]. Centre [254]. certain [320]. Changing [417, 33, 44]. characteristics [35, 51, 23, 25]. Chart [404, 351]. Checking [334, 269, 241]. Chinese [428]. Circulation [113, 131, 154, 167, 181, 196, 203, 214, 233, 223, 246, 258, 272, 282, 294]. Clarification [456]. Cleaning [152, 104, 174, 118]. Closing [263, 264]. CO [277]. CO-ALGOL [277]. code [256, 189]. Colloquium [263, 264].

Comment [151, 152, 66, 29, 68, 18, 47, 91, 94, 92, 93, 44, 91]. Commentaries [392, 408, 415]. commentary [355]. Comments [71, 315, 289, 436, 262, 16, 142, 187, 157, 254, 230, 118, 137, 122, 63, 140, 139, 138]. committee [9, 26]. Common [299]. Communications [60]. Companion [447]. Compilation [436, 451, 409, 442]. Compiled [374]. Compiler [434, 464, 85]. compilers [385]. Complete [1, 2, 294, 3, 4, 5, 6, 7]. Completer [329]. complex [76, 173]. compound [94]. Computer [237]. Computing [60, 254]. concept [186]. Concepts [207]. Concerning [208, 200, 288, 63, 50, 14, 13, 38, 21, 24, 30, 31, 69, 48]. conditional [45]. Conference [8, 70]. Consecutive [291]. Considerations [298]. considered [318]. constants [124, 159]. Construction [403, 35, 51, 85]. continued [160]. Contribution [212, 209, 230]. Contributions [148]. Contrived [399]. control [265]. convenient [147]. Conversational [426, 302, 313]. Conversion [381]. Copenhagen [9]. Copy [219]. Corrections [396]. Correspondence [337, 276, 297, 307, 84]. corresponding [66, 46]. Course [61]. Cripps [254]. Critique [266, 267]. Cyber [474]. cycle [445].

D. [120]. Data [300, 311, 420, 128]. Dealing [173]. December [20]. Decision [430]. Declaration [332]. Declarations [332, 76, 227]. Definition [287, 128, 63, 126, 30, 68, 189, 45]. Delimiters [43]. demonstrating [62]. denotations [330]. Describe [419]. Design [403]. designator [33, 44]. Development [212, 209, 230]. discontinuation [89]. Discussion [396, 75, 148]. Displays [299]. Distribution [52, 22, 73, 81, 87, 96]. Division [373, 368, 92]. Document [376]. Dr [138]. Draft [247, 234, 255, 273]. Dummy [291, 18]. during [75]. dynamic [67].

Early [476]. ECMALGOL [239]. editor [14, 13, 38, 21, 24, 316, 323, 342, 353, 362, 370, 372, 377, 379, 388, 405, 390, 407, 413, 421, 431, 433, 439, 441, 448, 453, 460, 455, 465, 470, 74, 99, 106, 114, 101, 108, 116, 133, 155, 168, 135, 158, 170, 182, 197, 204, 184, 199, 206, 224, 215, 226, 217, 235, 248, 259, 238, 250, 261, 274, 283, 286, 295, 305]. Editors [335]. efficiency [144]. efficient [147]. Effort [253]. element [42]. else [66, 46]. Enquiry [23]. Entry [430]. Environment [393]. Equivalence [331]. Errata [386, 387, 443, 442, 247, 352, 359, 369]. ERRATA-2 [352]. ERRATA-3 [359]. ERRATA-4 [369]. etc [25]. Europe [59]. evaluating [62]. Evaluation [49, 65, 93]. Example [62, 120, 33]. Exception [457, 458, 475]. Exception-Handling [458]. exchanges [160]. Executable [463]. Execution [430]. Experience [350]. Experiences [416]. Expressions [299, 333, 32, 65, 123, 132, 47, 93, 49]. extensible [322]. Extension [54, 237].

Facilities [436, 192]. Facility [367, 442, 409]. Factorial [399]. Failures [253]. Feasibility [411]. features [103]. few [149]. fewer [251]. File [220]. Files [417]. Final [338, 273]. Finite [418, 331]. first [82]. fixed [62]. FLACC [424]. Floating [423]. Floating-point [423]. Fontainebleau [319, 318, 317]. foregoing [140, 139]. form [14, 191]. Formal [287, 84, 419, 164, 262, 33, 321, 44]. Format [404]. Format-text [404]. Formatless [420]. Formula [57]. Formula-Translators [57]. Friendly [397]. function [33, 44]. functional [161]. Functionals [161]. Functions [423, 9]. Further [193, 201, 325].

General [326, 298]. Generalised [162]. Generalized [329]. generation [280]. given [66, 46]. Grammar [400]. Group [97, 102, 20, 36, 75]. Groups [332, 11].

H [398]. H. [157]. Handling [176, 201, 457, 458, 475, 119, 163, 147, 140, 138, 129]. Hardware [376, 395, 394, 418, 20]. Harware [384, 396]. High [451, 382, 345, 144]. High-Level [345]. Hoare [230, 210]. hope [150].

I/O [151, 122, 175, 207, 146, 239, 382]. I/O-Concepts [207]. I/O-proposal [175]. ICL [434]. ideas [160]. identifier [121, 164]. identifiers [162]. IFIP [376, 97, 102, 300, 301, 302, 303, 310, 311, 312, 313]. II [443]. imperfections [40]. Implementation [328, 474, 16, 53, 229, 447, 443, 298, 391, 68, 374]. Implementations [425, 424, 434, 444, 467, 473]. Implementing [468]. Implementors [360]. importance [62]. Improvements [325, 338, 301, 310, 315, 314, 318]. including [262]. Inclusion [76]. Indenter [437]. Independently [374]. Indirect [179, 177]. infinite [279, 280]. Initialisation [268, 80]. Input [420, 110, 55, 308, 163, 187]. Integer [221, 368, 124, 159, 173]. integers [31]. Integral [373]. Interactive [469]. Interface [303]. interpretation [358]. Invitations [160]. involving [280]. ISO [292, 82]. ISO-Representation [292]. Issue [1, 2, 3, 4, 5, 6, 7]. issues [38, 24]. Item [284]. items [260, 275, 296].

June [75].

Knuth [120].

labels [50, 31]. Language [402, 55, 410, 403, 411, 446, 234, 438, 265, 126, 143, 11, 125, 321, 322]. Languages [328, 451, 382, 302, 303, 313, 345, 218]. last [471]. Length [290]. Lengths [417]. Letter [390, 407, 455, 26, 40]. Letters [372, 379, 433, 441, 101, 108, 116, 135, 158, 170, 184, 199, 206, 226, 217, 238, 250, 261, 286]. Level [451, 382, 345]. levels [12]. Lexical [418]. Library [452, 464]. like [69]. Limited [430]. Line [417]. List [113, 131, 154, 167, 181, 196, 203, 214, 233, 223, 246, 294, 145, 258, 272, 282]. Living [446]. local [78]. Logic [419].

MABEL [410]. machine [12]. Machinery [60]. Macro [192].

Macro-Facilities [192]. magnetic [147]. Maintenance [71, 27, 39, 75, 301, 310, 64]. Mainz [20]. making [358]. Man [109, 136]. Manipulating [423]. Manipulation [429, 111, 242]. MARY [321, 332, 333]. masking [10]. Matching [419]. meaning [162]. Measurement [290]. Mechanism [173, 464, 458, 103]. Meeting [88, 317, 9, 20, 82]. Members [26, 36]. method [374]. minimum [326]. Minority [284, 273]. Misprints [34]. mixed [188, 117]. Modals [347]. Model [443]. modelling [103]. Modes [331, 279, 280]. Modified [456]. Modular [366, 451, 345]. Modules [436, 367, 409, 442]. Moscow [70]. Most [399]. MR93 [267, 254]. Multi [211]. Multi-programming [211]. Mutual [207].

N [138, 278]. N. [139]. NAG [452]. National [328]. Naur [143, 142]. Neighbourhood [397]. News [284, 260, 275, 296]. No [17, 304]. Non [142, 78]. Non-local [78]. Non-rectangular [142]. none [180]. Notational [231, 243]. Note [120, 339, 126, 179, 368, 67, 74, 82, 197, 86, 218, 194, 177, 335, 405, 421, 439, 248, 274]. Notes [210, 85, 224, 209, 316, 323, 342, 353, 362, 370, 377, 388, 413, 431, 448, 453, 460, 465, 470, 99, 106, 114, 133, 155, 168, 149, 182, 204, 215, 235, 259, 283, 295, 305]. Notice [14, 38, 21, 24, 287]. notions [69]. November [8]. number [190]. Numbers [423]. Numerical [393].

O [151, 122, 146, 239, 382]. O-Concepts [207]. O-proposal [175]. once [373]. One [175]. only [287]. Operating [397, 303, 313]. Operation [333]. operations [12, 10]. optimisation [269]. order [65, 62]. other [76, 244]. output [110, 308, 163, 187]. Overprinting [427]. own [80, 186, 67, 48].

P. [143]. Paper [137]. papers [86]. Parameterisation [393]. Parameters [41, 190, 77]. Parametrization [346, 364]. parentheses [29]. Paris [8]. Partial [346, 364]. Pattern [419]. Peculiarities [328]. philosophy [322]. Plea [211]. point [423]. primaries [62]. Problems [228]. Procedure [79, 121, 164, 50, 95, 78, 48, 129]. Procedures [239, 16, 53, 146]. processes [309]. Processing [220, 382, 300, 311, 270]. Program [399, 459, 327, 290, 403, 383, 30]. Programming [366, 451, 55, 410, 61, 462, 345, 411, 218, 211, 144]. Programs [411, 445, 280]. Progress [59, 35, 51]. Project [402]. Prolog [463]. Properties [411]. Proposal [426, 122, 221, 375, 210, 192, 367, 364, 457, 188, 475, 111, 151, 175, 163, 147, 140, 308, 139, 138]. Proposals [398, 436, 252, 319, 125, 172, 143, 157, 171]. Proposed [348, 54, 360, 194]. Pt [443]. Purpose [326].

quantities [78]. question [146, 309]. Questionnaire [360, 25, 83, 90, 365]. Quick [450]. quotes [127].

Random [417, 468]. reading [420]. real [173]. Recognition [330]. recognizer [384, 396]. Recommendations [289, 395]. Record [176, 201]. rectangular [142]. Recursive [16, 48]. Reference [450]. References [105, 112, 130, 180, 153, 166, 195, 202, 213, 232, 222, 245, 257, 271, 281, 293, 358]. referencing [309]. regarding [320]. Relations [32, 47, 117]. Release [425]. remaining [145]. Remark [63, 12, 10, 30, 31, 149, 43, 95, 72]. Remarks [435, 143, 401, 212, 178, 193, 255]. removing [40]. replicating [459, 445]. Replies [90, 365]. Reply [28, 140, 139, 159]. Report [284, 376, 318, 317, 325, 338, 387, 9, 247, 304, 300, 301, 302, 303, 273, 234, 438, 255, 40, 355, 319, 352, 359, 369, 392, 408, 415, 34, 37, 98, 341]. Representation [221, 384, 396, 376, 395, 394, 418, 292]. representations [320]. Reprinted [60]. reproducing [327]. review [349]. Revised [355, 352, 359, 369, 387, 438, 404, 341, 392, 408, 415]. Revision [360, 319]. Revisions [71]. Rome [88]. Rounding [265]. Routine [374, 330]. routine-denotations [330]. Routines [381]. Row [299]. Row-Displays [299]. Rule [219]. run [241]. run-time [241]. Rutishauser [157].

SBTNon [11]. SBTNon-ALGOL [11]. Schema [420]. Scheme [56]. Science [462]. Scope [334, 50]. sections [319]. Self [459, 327, 445]. Self-replicating [459, 445]. Self-reproducing [327]. Semantics [463]. Semicolons [291]. Separate [436, 409, 442]. Set [242, 239]. Sets [244, 262]. shifting [10]. signed [124, 159]. SIL [350]. Simple [375, 103, 308]. Simulation [340, 357, 356, 228]. Sliced [299]. Software [393]. Some [122, 178, 255, 385, 298, 321, 171, 139, 138, 103, 476]. Specification [402, 364]. specifications [191]. spot [165]. Spots [394, 145]. Standard [384, 396, 376, 394, 418, 9]. Standardise [221]. State [418]. statement [185, 174, 18, 95]. Statements [219, 291, 240, 91, 94, 45]. Static [334]. step [42]. step-until [42]. Stockholm [11]. Storage [56, 15, 12]. String [119, 207, 127, 111, 140, 138, 129]. string-handling [140, 138]. Strings [77]. Structure [57]. Sub [299, 9]. sub-committee [9]. Sub-Expressions [299]. Subcommittee [344, 300, 301, 302, 303, 310, 311, 312, 313]. Sublanguage [326, 348, 361]. Sublanguages [312]. Subscript [269]. Successes [253]. successor [194]. such [150]. suggested [315]. Suggestion [64, 65, 50, 69, 45]. Suggestions [314, 320, 11, 13, 70, 48]. Summary [75, 365]. Superfluous [43]. Superlanguage [398]. Supplement [386, 17, 98]. Supplements [13, 52, 22, 73, 81, 87, 96]. Support [207, 344]. Survey [444, 467, 473]. Symbols [221]. syntactic [189]. Syntax [404, 375, 288, 351, 383, 58, 243]. System [55, 397, 419, 308, 144]. Systems [451, 303, 313].

Tables [430]. tapes [147]. Teachability [382]. Teaching [208, 134, 200]. tears [251]. TESLA [391]. test [120]. Text [270, 404, 273]. Texts [374]. there [150]. Thoughts [412, 201]. time [241]. Timing [240]. Token [384, 396]. Token-recognizer [384, 396]. tokens [256]. transformations [161]. Translation [428]. Translations [37]. translator [35, 51, 23, 25].

Translators [57, 69]. Transportable [393]. Transput [426, 435, 443, 300, 311, 381, 416, 319]. treatment [262]. Trouble [394, 145, 165]. two [12, 86]. Type [95, 241, 121, 84, 188]. types [173, 262, 244, 128, 117].

U.S. [75]. unit [375]. until [42]. Use [419, 350, 29]. Usefulness [411].

value [33, 44, 41]. Values [332]. variable [33, 44]. variables [268, 80]. vary [190]. versus [68]. Viable [467, 473, 444]. View [356]. Visibility [382].

WG [344]. WG2.1 [317, 300, 301, 302, 303, 310, 311, 312, 313]. Wijngaarden [149, 137]. Wirth [230, 139, 138]. Withdrawal [227]. within [110, 111]. Word [263, 264]. Working [70, 97, 102].

X [186, 171, 194, 172, 128].

zero [92]. Zurich [264].

References

Naur:1959:CIa

[1] Peter Naur. Complete issue. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (1):1–6, March 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:CIb

[2] Peter Naur. Complete issue. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (2):1–7, May 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:CIc

[3] Peter Naur. (Complete issue). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (3):1–6, June 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:CId

[4] Peter Naur. (Complete issue). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (4):1–7, August 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:CIe

[5] Peter Naur. (Complete issue). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (5):1–9, August 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:CIf

[6] Peter Naur. (Complete issue). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (6):1–2, October 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:CIg

[7] Peter Naur. (Complete issue). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (7):1–21, November 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:ACP

[8] Peter Naur. ALGOL — conference in Paris, 12–14 November 1959. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (8):1–7, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:RSC

[9] Peter Naur. Report of the sub-committee of the meeting in Copenhagen on Standard Functions. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (8):7–8, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Garwick:1959:RSM

[10] Jan V. Garwick. Remark on shifting and masking operations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (8):9, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:SAG

[11] Peter Naur. Suggestions for the ALGOL groups in Stockholm @SBTNon-ALGOL language. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (8):9, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Garwick:1959:ROM

[12] Jan V. Garwick. Remark on operations for a machine with two levels of storage. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (8):10, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:SEC

[13] Peter Naur. Suggestions from the editor concerning ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (8):10–11, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Naur:1959:NEC

[14] Peter Naur. Notice from the editor concerning the form of the ALGOL Bulletin. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (8):11, December 1959. CODEN ALGOBG. ISSN 0084-6198.

Sattley:1960:ASA

[15] K. Sattley and P. Z. Ingerman. The allocation of storage for arrays in Algol 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 13.1):1–15, November 1960. CODEN ALGOBG. ISSN 0084-6198.

Irons:1960:CIR

[16] E. T. Irons and W. Feurzeig. Comments on the implementation of recursive procedures and blocks in Algol-60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 13.2):1–15, November 1960. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1960:ASN

[17] Peter Naur. ALGOL supplement no. 5. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 5):1–29, November 1960. CODEN AL-GOBG. ISSN 0084-6198.

Rutishauser:1960:CDS

[18] Heinz Rutishauser. Comment [on the dummy statement]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (5):1, September 28, 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:A

[19] Peter Naur. ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (9):1, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:AHG

[20] Peter Naur. ALCOR hardware group meeting in Mainz, December 17–19, 1959. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (9):1–2, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:NECb

[21] Peter Naur. Notice from the editor concerning algorithms in CACM. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (9):2, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:DABb

[22] Peter Naur. Distribution of ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (9):3, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:EAT

[23] Peter Naur. Enquiry into ALGOL translator characteristics. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (9):3, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:NECc

[24] Peter Naur. Notice from the editor concerning back issues. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (9):3, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:QAT

[25] Peter Naur. Questionnaire on ALGOL translator characteristics, etc. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (9):4, March 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:LMA

[26] Peter Naur. Letter to the members of the ALGOL 60 Committee. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (10):1–2, October 17, 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:AMa

[27] Peter Naur. ALGOL 60 maintenance. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):1–10, October 1960. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1960:R

[28] Heinz Rutishauser. Reply to AB 10.1.3. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):2–3, October 17, 1960. CODEN ALGOBG. ISSN 0084-6198. See [26].

Kerner:1960:CUP

[29] J. Kerner. Comment of the use of parentheses. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):10, October 1960. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1960:RCD

[30] Peter Naur. Remark concerning the definition of a program. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):11, October 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:RCI

[31] Peter Naur and I. Dahlstrand. Remark concerning integers as labels. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):11, October 1960. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1960:ER

[32] Heinz Rutishauser. Expressions in relations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):11, October 17, 1960. CODEN

ALGOBG. ISSN 0084-6198. URL https://dl.acm.org/doi/10.5555/
1060892.1060896.

Naur:1960:EFD

[33] Peter Naur. An example of a function designator changing the value of a formal variable. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):12–13, October 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:MAR

[34] Peter Naur. Misprints in the ALGOL 60 Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):13, October 1960. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1960:ATCa

[35] Peter Naur. ALGOL translator characteristics and the progress in translator construction. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):14–16, October 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:MAG

[36] Peter Naur. Members of the ALCOR group. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):16, October 1960. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1960:TAR

[37] Peter Naur. Translations of the ALGOL 60 Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):16, October 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:NECa

[38] Peter Naur. Notice from the editor concerning back issues. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (10):17, October 1960. CO-DEN ALGOBG. ISSN 0084-6198.

Naur:1960:AMb

[39] Peter Naur. ALGOL 60 maintenance. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):1–4, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1960:LRI

[40] H. Rutishauser. Letter [removing imperfections from the ALGOL report]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):2, December 23, 1960. CODEN ALGOBG. ISSN 0084-6198. URL https://dl.

acm.org/doi/10.5555/1060905.1060912.

Woodger:1960:PCV

[41] M. Woodger. Parameters called by value. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):5, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1960:SUE

[42] M. Woodger. The step-until element. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):5, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1960:RSD

[43] H. Rutishauser. Remark on superfluous delimiters. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):6, December 23, 1960. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1960:CFD

[44] M. Woodger. A comment on a function designator changing the value of a formal variable. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):6, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1960:SDC

[45] M. Woodger. Suggestion on the definition of conditional statements. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (11):6, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:TCG

[46] Peter Naur. The then corresponding to a given else. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):7, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1960:CER

[47] H. Rutishauser. Comment on expressions in relations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):7, December 23, 1960. CO-DEN ALGOBG. ISSN 0084-6198. URL https://dl.acm.org/doi/10.

5555/1060905.1060912.

vanderPoel:1960:SCO

[48] W. L. van der Poel. Suggestions concerning own in recursive procedure. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):7, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1960:EE

[49] M. Woodger. Evaluation of expressions. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):8–9, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Mayoh:1960:SCS

[50] B. H. Mayoh. Suggestion concerning the scope of labels in procedure bodies. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):9, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:ATCb

[51] Peter Naur. ALGOL translator characteristics and the progress in translator construction. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):10, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Naur:1960:DABa

[52] Peter Naur. Distribution of ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (11):10, December 1960. CO-DEN ALGOBG. ISSN 0084-6198.

Jensen:1961:IAP

[53] J. Jensen and Peter Naur. An implementation of Algol 60 procedures. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 11):38– 47, January 1961. CODEN ALGOBG. ISSN 0084-6198. Preprint from Nordisk Tidskrift for Informations-Behandling, Volume 1, No 1, 1961.

Hockney:1961:PEA

[54] R. W. Hockney. A proposed extension to Algol 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 12):1–12, June 1961. CO-DEN ALGOBG. ISSN 0084-6198.

Ershov:1961:ILS

[55] P. Ershov, G. I. Koschuhin, and U. M. Voloshin. The input language system of automatic programming. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 14):1–9, January 1961. CODEN ALGOBG. ISSN 0084-6198.

Jensen:1961:SAS

[56] J. Jensen, P. Mondrup, and Peter Naur. A storage allocation scheme for ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 15):1–12, October 1961. CODEN ALGOBG. ISSN 0084-6198.

Lucas:1961:SFT

[57] P. Lucas. The structure of formula-translators. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 16):1–27, September 1961. CODEN ALGOBG. ISSN 0084-6198.

Youden:1961:AAS

[58] William W. Youden. An analysis of the ALGOL syntax. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 17):1–13, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:PAE

[59] Peter Naur. Progress of ALGOL in Europe. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 18):1–6, October 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:RCA

[60] Peter Naur. Reprinted from the Communications of the Association for Computing Machinery. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 6):1–6, January 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:CAP

[61] Peter Naur. A course of Algol 60 programming. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 9):1–38, January 1961. CODEN ALGOBG. ISSN 0084-6198.

Dijkstra:1961:EDI

[62] E. W. Dijkstra. Example demonstrating the importance of evaluating the primaries in a fixed order. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):1, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Dijkstra:1961:RCD

[63] E. W. Dijkstra. Remark concerning the definition of comments. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):1, April 1961. CO-DEN ALGOBG. ISSN 0084-6198.

Dahlstrand:1961:SAM

[64] I. Dahlstrand. Suggestion on ALGOL 60 maintenance. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):2, April 1961. CODEN AL-GOBG. ISSN 0084-6198.

Dahlstrand:1961:SOE

[65] I. Dahlstrand. Suggestion on the order of evaluation of expressions. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (12):2, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Dahlstrand:1961:CTC

[66] I. Dahlstrand. Comment on the then corresponding to a given else. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (12):3, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:NDO

[67] Peter Naur. Note on dynamic own arrays. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):3, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:CDV

[68] Peter Naur. Comment on the definition versus implementation of AL-GOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):4, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:SCN

[69] Peter Naur. Suggestion concerning the notions of ALGOL and ALGOLlike translators. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):4, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:SMW

[70] Peter Naur. Suggestions from the Moscow Working Conference on AL-GOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):4–6, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Bauer:1961:CAM

[71] F. L. Bauer and K. Samelson. Comments on ALGOL 60 maintenance and revisions. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12): 7–8, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Ershov:1961:R

[72] A. P. Ershov. A remark. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):8, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:DABa

[73] Peter Naur. Distribution of ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):9, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:NE

[74] Peter Naur. Note from the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (12):9, April 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:SDU

[75] Peter Naur. Summary of the discussion in the U.S. ALGOL 60 Maintenance Group during August 1960 to June 1961. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):1–11, August 1961. CODEN AL-GOBG. ISSN 0084-6198.

Hockney:1961:ICO

[76] R. W. Hockney. Inclusion of complex and other declarations in ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):11, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1961:SAP

[77] M. Woodger. Strings as actual parameters. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):11, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1961:NLQ

[78] M. Woodger. Non-local quantities of procedure bodies. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):12, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1961:PBB

[79] M. Woodger. Procedure bodies as blocks. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):12, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Mayoh:1961:IOV

[80] B. H. Mayoh. Initialisation of own variables. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):12–13, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1961:DABb

[81] Peter Naur. Distribution of ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):13, August 1961. CO-DEN ALGOBG. ISSN 0084-6198.

Naur:1961:NFM

[82] Peter Naur. Note on the first meeting of ISO. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (13):13, August 1961. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:Q

[83] Peter Naur. The questionnaire. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (14):1–14, January 1962. CODEN ALGOBG. ISSN 0084-6198.

Nagao:1962:AFC

[84] M. Nagao. Actual-formal correspondence in type. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (14):14, January 1962. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1962:NCC

[85] Peter Naur. Notes on compiler construction. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (14):14–15, January 1962. CODEN ALGOBG. ISSN 0084-6198.

Sammet:1962:NTP

[86] Jean E. Sammet. Note on two papers. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (14):15, January 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:DABa

[87] Peter Naur. Distribution of ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (14):16, January 1962. CO-DEN ALGOBG. ISSN 0084-6198.

Naur:1962:AMR

[88] Peter Naur. The ALGOL meeting in Rome. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):1–2, June 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:DABb

[89] Peter Naur. The discontinuation of the ALGOL Bulletin. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):2–3, June 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:RAQ

[90] Peter Naur. The replies to the AB14 questionnaire. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):3–51, June 1962. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1962:CCSa

[91] Heinz Rutishauser. Comment [on comment statements]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):17, June 16, 1962. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1962:CDZ

[92] Heinz Rutishauser. Comment [on division by zero]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):20, June 16, 1962. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1962:CEE

[93] Heinz Rutishauser. Comment [on evaluation of expressions]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):21, June 16, 1962. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1962:CCSb

[94] Heinz Rutishauser. Comment [on compound statements]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):22, June 16, 1962. CO-DEN ALGOBG. ISSN 0084-6198.

Rutishauser:1962:RTP

[95] Heinz Rutishauser. Remark [type procedure called by procedure statement]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):34, June 16, 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:DABc

[96] Peter Naur. Distribution of ALGOL Bulletin Supplements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):52, June 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:IWG

[97] Peter Naur. The IFIP working group on ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):52, June 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1962:SAR

[98] Peter Naur. Supplement to the ALGOL 60 Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (15):53–57, June 1962. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:ENa

[99] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):1–2, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:Aa

[100] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):3, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:LEa

[101] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):3, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:IWG

[102] Peter Naur. IFIP working group on ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):4–13, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Dijkstra:1964:SMM

[103] E. W. Dijkstra. A simple mechanism modelling some features of ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):14–23, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Duncan:1964:CA

[104] F. G. Duncan and A. van Wijngaarden. Cleaning up ALGOL60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):24–32, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:Ra

[105] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (16):33–34, May 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:ENb

[106] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):1, July 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:Ab

[107] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):2–4, July 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:LEb

[108] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):5–10, July 1964. CODEN ALGOBG. ISSN 0084-

Knuth:1964:MB

[109] Donald E. Knuth. Man or boy? ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):7, January 1964. CODEN ALGOBG. ISSN 0084-6198.

Brokate:1964:IOW

[110] K. G. Brokate. Input and output within ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):11–12, July 1964. CODEN AL-GOBG. ISSN 0084-6198.

Wirth:1964:PSM

[111] N. Wirth. A proposal on string manipulation within ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):13–17, July 1964. CODEN ALGOBG. ISSN 0084-6198. See comments [138] and reply [140].

Naur:1964:Rb

[112] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):18, July 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:CLa

[113] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17):19–29, July 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:ENc

[114] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):1, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:Ac

[115] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):2–3, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:LEc

[116] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):4–9, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Thacher:1964:MTR

[117] H. C. Thacher. On mixed types in relations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):10, October 1964. CODEN AL-GOBG. ISSN 0084-6198.

Thacher:1964:CCA

[118] H. C. Thacher. Comments on “Cleaning up ALGOL60”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):10–11, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Cook:1964:SHA

[119] R. L. Cook. String handling in ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):12–13, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Bekic:1964:NTE

[120] H. Bekic. Note on a test example by D. E. Knuth. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):13, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Bekic:1964:ATP

[121] H. Bekic. The assignment to a type procedure identifier. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):14, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Bekic:1964:SCA

[122] H. Bekic. Some comments on the ACM I/O proposal. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):16–19, October 1964. CO-DEN ALGOBG. ISSN 0084-6198.

Hoare:1964:CEa

[123] C. A. R. Hoare. Case expressions. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):20–22, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Petrone:1964:ISC

[124] L. Petrone and C. E. Vandoni. Integer and signed constants in ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):23–25, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:PNL

[125] P. Naur. Proposals for a new language. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):26–43, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Dijkstra:1964:NLD

[126] E. W. Dijkstra. Note on language definition. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):44, October 1964. CODEN AL-GOBG. ISSN 0084-6198.

Randell:1964:SQ

[127] B. Randell. String quotes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):44, October 1964. CODEN ALGOBG. ISSN 0084-6198.

McCarthy:1964:DND

[128] J. McCarthy. Definition of new data types in ALGOL x. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):45, October 1964. CODEN ALGOBG. ISSN 0084-6198.

vanderPoel:1964:PSH

[129] W. L. van der Poel. A procedure for string handling. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):47–48, October 1964. CO-DEN ALGOBG. ISSN 0084-6198.

Naur:1964:Rc

[130] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):49–50, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Naur:1964:CLb

[131] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (18):51–53, October 1964. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1964:CEb

[132] C. A. R. Hoare. Case expressions. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):20–22, ???? 1964. CODEN ALGOBG. ISSN

0084-6198. URL https://dl.acm.org/doi/10.5555/1060978.1060988.

Naur:1965:ENa

[133] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):1, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Knuth:1965:TA

[134] Donald E. Knuth. Teaching ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):4–6, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:LEa

[135] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):4–10, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Knuth:1965:MB

[136] Donald E. Knuth. Man or boy? ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):8–9, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Wirth:1965:CPV

[137] N. Wirth. Comments on a paper by A. van Wijngaarden. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):11–12, January 1965. CODEN ALGOBG. ISSN 0084-6198.

vandeLaarschot:1965:SCS

[138] P. J. J. van de Laarschot and J. Nederkoorn. Some comments on the string-handling proposal of Dr N. Wirth. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):12–13, January 1965. CODEN AL-

GOBG. ISSN 0084-6198. URL https://dl.acm.org/doi/pdf/10.5555/

1060998.1061002. See [111] and reply [140].

vandeLaarschot:1965:SCP

[139] P. J. J. van de Laarschot and J. Nederkoorn. Some comments on a proposal by N. Wirth; N. Wirth: Reply to the foregoing comments. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):12–14, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Wirth:1965:RFC

[140] N. Wirth. Reply to the foregoing comments [on the string-handling proposal]. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):14, January 1965. CODEN ALGOBG. ISSN 0084-6198. URL https://dl.

acm.org/doi/pdf/10.5555/1060998.1061002. See [111, 138].

Ingerman:1965:BA

[141] P. Z. Ingerman. Blemishes of ALGOL60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):14–15, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Landin:1965:CNN

[142] P. J. Landin. Comments on Naur’s “Non-rectangular arrays”. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (19):15–17, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Garwick:1965:RPN

[143] J. V. Garwick. Remarks on P. Naur’s new language proposals. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (19):17–18, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Ershov:1965:AAP

[144] A. P. Ershov. ALPHA — an automatic programming system of high efficiency. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19): 19–27, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Knuth:1965:LRT

[145] Donald E. Knuth. A list of the remaining trouble spots in ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):29–38, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Garwick:1965:QP

[146] J. V. Garwick. The question of I/O procedures. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):39–41, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Seegmuller:1965:PEC

[147] G. Seegmuller. A proposal for efficient and convenient handling of magnetic tapes in ALGOL60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):42–43, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1965:CDA

[148] H. Rutishauser. Contributions to the discussion of ALGOL 6x. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (19):44–50, January 1965. CODEN ALGOBG. ISSN 0084-6198. URL https://dl.acm.org/

doi/10.5555/1060998.1061010. See comments [157].

Naur:1965:FNV

[149] P. Naur. A few notes; A. van Wijngaarden: Remark on AB19.3.11.4. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (19):51–55, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Engstrom:1965:TSA

[150] G. Engstrom. “But there are such animals — I hope”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):55–56, January 1965. CO-DEN ALGOBG. ISSN 0084-6198.

Ageev:1965:CAP

[151] M. I. Ageev. Comment on the ACM I/O proposal. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):57, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Bekic:1965:CCA

[152] H. Bekic. Comment on “Cleaning up ALGOL60”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):58, January 1965. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1965:Ra

[153] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):59, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:CLa

[154] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):60–63, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:ENb

[155] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):1–2, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:Aa

[156] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):3–4, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Myrhaug:1965:CPH

[157] Björn Myrhaug. Comments on the proposals of H. Rutishauser, AB19.3.10. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20): 5–6, July 1965. CODEN ALGOBG. ISSN 0084-6198. See [148].

Naur:1965:LEb

[158] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):5–20, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Petrone:1965:ISC

[159] L. Petrone. Integer and signed constants — a reply. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):21–22, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:ICE

[160] Peter Naur. Invitations to continued exchanges of ideas. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):23–26, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Samelson:1965:FFT

[161] K. Samelson. Functionals and functional transformations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):27–28, July 1965. CO-DEN ALGOBG. ISSN 0084-6198.

Bekic:1965:MIA

[162] H. Bekic. The meaning of identifiers in ALGOL and in “Generalised ALGOL”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20): 29–30, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Hoffman:1965:PIO

[163] H. J. Hoffman. A proposal for Input–Output handling in ALGOL60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):31–45, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Bekic:1965:AFP

[164] H. Bekic. Assignment to formal procedure identifier. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):45–46, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Medema:1965:ATS

[165] P. Medema. Another trouble spot in ALGOL60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):47–48, July 1965. CODEN AL-GOBG. ISSN 0084-6198.

Naur:1965:Rb

[166] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):49, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:CLb

[167] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (20):50–55, July 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:ENc

[168] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):1, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:Ab

[169] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):2–4, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:LEc

[170] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):5–6, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Seegmuller:1965:SPA

[171] G. Seegmuller. Some proposals for ALGOL X. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):7–22, November 1965. CODEN ALGOBG. ISSN 0084-6198.

vonPeschke:1965:PAX

[172] Joachim von Peschke. Proposals for ALGOL X. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):23, November 1965. CODEN AL-GOBG. ISSN 0084-6198.

Dijkstra:1965:MDT

[173] E. W. Dijkstra. A mechanism for dealing with the types integer, real and complex. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):24–31, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1965:CS

[174] C. A. R. Hoare. Cleaning up the for statement. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):32–35, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Durchholz:1965:OMP

[175] R. Durchholz. One more I/O-proposal for ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):36–38, November 1965. CO-DEN ALGOBG. ISSN 0084-6198.

Hoare:1965:RH

[176] C. A. R. Hoare. Record handling. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):39–69, November 1965. CODEN ALGOBG. ISSN

0084-6198. URL https://dl.acm.org/doi/10.5555/1061032.1061041.

Hoare:1965:NIAa

[177] C. A. R. Hoare. A note on indirect addressing. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):63–65, November 1965. CODEN

ALGOBG. ISSN 0084-6198. URL https://archive.computerhistory.
org/resources/text/algol/algol_bulletin/A21/P38.HTM.

Hill:1965:SRA

[178] I. D. Hill. Some remarks on ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):70–74, November 1965. CODEN AL-GOBG. ISSN 0084-6198.

Hoare:1965:NIAb

[179] C. A. R. Hoare. A note on indirect addressing. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):75–77, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:RN

[180] Peter Naur. References (none). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):78, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1965:CLc

[181] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (21):79–83, November 1965. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:ENa

[182] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):1–2, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:Aa

[183] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):3, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:LEa

[184] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):4–5, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Galler:1966:SA

[185] B. A. Galler. The for statement in ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):6–7, February 1966. CODEN AL-GOBG. ISSN 0084-6198.

Merner:1966:OCA

[186] J. N. Merner. own concept and ALGOL X. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):7–9, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Merner:1966:CIO

[187] J. N. Merner. Comments on Input–Output. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):9–10, February 1966. CODEN AL-GOBG. ISSN 0084-6198.

Newman:1966:PTM

[188] G. B. Newman. Proposal for type mixed. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):11, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Shogolev:1966:CSD

[189] E. A. Shogolev. On code syntactic definition. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):11–12, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:ANP

[190] Peter Naur. Allowing number of parameters to vary. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):13, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:FS

[191] Peter Naur. The form of specifications. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):14, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Leroy:1966:PMF

[192] H. Leroy. A proposal for macro-facilities in ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):15–26, February 1966. CO-DEN ALGOBG. ISSN 0084-6198.

Hill:1966:FRA

[193] I. D. Hill. Further remarks on ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):27, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1966:AXN

[194] M. Woodger. ALGOL X, note on the proposed successor to ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):28–33, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:Ra

[195] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):34, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:CLa

[196] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (22):35–38, February 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:ENb

[197] Peter Naur. Editor’s note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):1, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:Ab

[198] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):2, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:LEb

[199] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):3–4, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Lennie:1966:CTA

[200] John Lennie. Concerning “Teaching ALGOL 60”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):4, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1966:FTR

[201] C. A. R. Hoare. Further thoughts on record handling AB21.3.6. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):5–11, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:Rb

[202] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):12, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:CLb

[203] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (23):13–15, May 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:ENc

[204] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):1, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:Ac

[205] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):2–4, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:LEc

[206] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):5–8, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Durchholz:1966:MSS

[207] R. Durchholz. Mutual support of string and I/O-concepts. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):9–11, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Hill:1966:CTA

[208] I. D. Hill. Concerning “Teaching ALGOL 60”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):12, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Wirth:1966:ANC

[209] N. Wirth. Additional notes on “A Contribution to the Development of ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24): 13–17, September 1966. CODEN ALGOBG. ISSN 0084-6198. See comments [210].

Kral:1966:NHP

[210] Jaroslav Král. A notes on Hoare Proposal AB21.3.6. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):17–25, September 1966. CO-DEN ALGOBG. ISSN 0084-6198. See [209].

Dahl:1966:PMP

[211] O. J. Dahl. A plea for multi-programming. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):26–27, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Hill:1966:RCD

[212] I. D. Hill. Remarks on “Contribution to the Development of ALGOL”. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (24):28–32, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:Rc

[213] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):33, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1966:CLc

[214] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):34–37, September 1966. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:ENb

[215] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):1–3, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:Ab

[216] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):3–5, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:LEb

[217] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):5–7, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Wood:1967:NPL

[218] D. Wood. A note on programming languages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):7–10, March 1967. CODEN AL-GOBG. ISSN 0084-6198.

Hill:1967:SCR

[219] I. D. Hill. for statements and the copy rule. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):11–12, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1967:FP

[220] C. A. R. Hoare. File processing. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):12–19, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Bell:1967:PSI

[221] A. G. Bell, F. R. A. Hopgood, and M. Woodger. Proposal to standardise an integer representation for ALGOL basic symbols. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):19–24, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:Rb

[222] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):25, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:CLb

[223] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (25):26–30, March 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:ENa

[224] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):1, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:Aa

[225] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):2–4, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:LEa

[226] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):5–14, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Loof:1967:WD

[227] H. V. Loof. Withdrawal of declarations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):18–19, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Kral:1967:PS

[228] J. Kral. Problems of simulation. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):20–21, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Kral:1967:IA

[229] J. Kral. Implementation of AB24.3.4. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):22, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Simauti:1967:CCD

[230] T. Simauti et al. Comments on Contribution to the development of AL-GOL (Wirth and Hoare). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):23–27, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Duncan:1967:NA

[231] F. G. Duncan. Notational abbreviations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):28–42, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:Ra

[232] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):43, July 1967. CODEN ALGOBG. ISSN 0084-6198.

Naur:1967:CLa

[233] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (26):44–48, July 1967. CODEN ALGOBG. ISSN 0084-6198.

vanWijngaarden:1968:DRA

[234] A. van Wijngaarden, B. J. Mailloux, J. Peck, and C. H. A. Koster. Draft report on the algorithmic language ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 26):1–84, March 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:ENa

[235] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):1, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:Aa

[236] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):2–5, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Wirth:1968:ANC

[237] Niklaus Wirth. An extension of ALGOL on the 360 computer. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):4–5, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:LEa

[238] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):6–12, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Hekimi:1968:SPE

[239] D. Hekimi. A set of I/O procedures for ECMALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):13–23, February 1968. CO-DEN ALGOBG. ISSN 0084-6198.

Wichmann:1968:TAS

[240] B. A. Wichmann. Timing ALGOL statements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):24–26, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Scowen:1968:TCR

[241] R. S. Scowen. Type checking at run-time. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):27–28, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1968:SM

[242] C. A. R. Hoare. Set manipulation. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):29–37, February 1968. CODEN ALGOBG. ISSN

0084-6198. URL https://dl.acm.org/doi/10.5555/1061098.1061105.

Gries:1968:NAA

[243] D. Gries. Notational abbreviations applied to the syntax of ALGOL. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (27):38–40, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Laski:1968:SOT

[244] J. G. Laski. Sets and other types. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):41–48, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:Ra

[245] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):49, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:CLa

[246] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):50–54, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:EDR

[247] Peter Naur. Errata to the draft report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (27):55–64, February 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:ENb

[248] Peter Naur. Editor’s note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):1, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:Ab

[249] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):2–3, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:LEb

[250] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):4–8, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1968:AFT

[251] C. H. Lindsey. ALGOL68 with fewer tears. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):9–49, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1968:PAA

[252] C. H. Lindsey. Proposals for amendment of ALGOL68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):50–57, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:SFA

[253] Peter Naur. Successes and failures of the ALGOL effort. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):58–62, July 1968. CO-

DEN ALGOBG. ISSN 0084-6198. URL http://doi.acm.org/10.1145/

1061112.1061118. Talk given at the colloquium “Ten Years of Algol”, Eidgenössische Technische Hochschule, Zürich, May 31, 1968.

Naur:1968:CCC

[254] Peter Naur. Cripps Computing Centre: Comments on MR93. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):63–64, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Hill:1968:SRD

[255] I. D. Hill. Some remarks on the draft report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):65–69, July 1968. CODEN AL-GOBG. ISSN 0084-6198.

Baecker:1968:ACA

[256] H. D. Baecker. ASERC, a code for ALGOL68 basic tokens. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):70–75, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:Rb

[257] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):76, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:CLb

[258] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (28):77–82, July 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:ENc

[259] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):1, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:NI

[260] Peter Naur. News items. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):2, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:LEc

[261] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):3–11, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Igarashi:1968:CFT

[262] S. Igarashi. Comments on the formal treatment of types including sets. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):12– 15, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Wirth:1968:ACC

[263] Niklaus Wirth. ALGOL colloquium — closing word. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):16–19, May 31, 1968. CODEN ALGOBG. ISSN 0084-6198. URL http://archive.

computerhistory.org/resources/text/algol/ACM_Algol_bulletin/
1061166/p16-wirth.pdf.

Wirth:1968:CWZ

[264] N. Wirth. Closing word at Zurich Colloquium. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):16–19, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Apostolatos:1968:RCA

[265] N. Apostolatos, H. Christ, H. Santo, and H. Wippermann. Rounding control and the algorithmic language ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):20–26, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1968:CA

[266] C. A. R. Hoare. Critique of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):27–29, November 1968. CODEN AL-

GOBG. ISSN 0084-6198. URL https://archive.computerhistory.
org/resources/text/algol/algol_bulletin/A29/P34.HTM.

Hoare:1968:CM

[267] C. A. R. Hoare. Critique of MR93. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):27–29, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Hoare:1968:IV

[268] C. A. R. Hoare. Initialisation of variables. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):30–32, November 1968. CODEN AL-GOBG. ISSN 0084-6198.

Hoare:1968:SOC

[269] C. A. R. Hoare. Subscript optimisation and checking. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):33–44, November 1968. CO-DEN ALGOBG. ISSN 0084-6198.

Hoare:1968:TP

[270] C. A. R. Hoare. Text processing. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):45–60, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:Rc

[271] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):61, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Naur:1968:CLc

[272] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):62–68, November 1968. CODEN ALGOBG. ISSN 0084-6198.

Woodger:1969:MRA

[273] Mike Woodger, Władisław Marek Turski, Edsger W. Dijkstra, C. A. R. Hoare, Brian Randell, Jan V. Garwick, Fraser G. Duncan, and Gerhard Seegmueller. Minority report on Algol 68. The draft and the final text. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (??):??, January

17, 1969. CODEN ALGOBG. URL http://ershov.iis.nsk.su/en/
node/805785.

Naur:1969:EN

[274] Peter Naur. Editor’s note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):1, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Naur:1969:NI

[275] Peter Naur. News items. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):2–3, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Naur:1969:C

[276] Peter Naur. Correspondence. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):4–16, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Roberts:1969:CA

[277] J. D. Roberts. CO-ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):17–37, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Igarashi:1969:A

[278] S. Igarashi, T. lwamura, K. Sakuma, T. Simauti, T. Simuzu, S. Takasu, E. Wada, and N. Yoneda. ALGOL N. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):38–85, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Koster:1969:IM

[279] C. H. A. Koster. On infinite modes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):86–89, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Meertens:1969:GAP

[280] L. Meertens. On the generation of ALGOL68 programs involving infinite modes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30): 90–92, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Naur:1969:R

[281] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):93, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Naur:1969:CL

[282] Peter Naur. Circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (30):94–97, February 1969. CODEN ALGOBG. ISSN 0084-6198.

Naur:1970:EN

[283] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):1–6, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Dijkstra:1970:NIM

[284] Edsger W. Dijkstra, Fraser G. Duncan, Jan V. Garwick, C. A. R. Hoare, Brian Randell, G. Seegmueller, Władisław M. Turski, and Michael Woodger. News item — minority report. ALGOL Bulletin (Amsterdam:

Mathematisch Centrum), (31):7, March 1970. CODEN ALGOBG. ISSN 0084-6198. Item AB31.1.1.

Naur:1970:A

[285] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):7–9, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Naur:1970:LE

[286] Peter Naur. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):9–15, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Lauer:1970:FDA

[287] P. E. Lauer. Formal definition of ALGOL 60 (notice only). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):16, March 1970. CO-DEN ALGOBG. ISSN 0084-6198.

Pair:1970:CSA

[288] C. Pair. Concerning the syntax of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):16–27, March 1970. CODEN AL-GOBG. ISSN 0084-6198.

Bowlden:1970:ACR

[289] H. J. Bowlden. ALGOL 68 — comments and recommendations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):28–32, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Palme:1970:AMP

[290] J. Palme. “Atoms“ as a measurement of program length. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):33–34, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Nievergelt:1970:DSC

[291] J. Nievergelt. Dummy statements and consecutive semicolons in ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):35–36, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1970:IRA

[292] C. H. Lindsey. An ISO-representation for ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):37–60, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Naur:1970:R

[293] Peter Naur. References. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):61–62, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Naur:1970:NCC

[294] Peter Naur. The new complete circulation list. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (31):62, March 1970. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:EN

[295] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):1–2, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:NI

[296] Peter Naur. News items. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):3–4, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:C

[297] Peter Naur. Correspondence. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):5–17, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Molnar:1971:SGC

[298] G. Molnar. Some general considerations of implementation. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):18–21, May 1971. CO-DEN ALGOBG. ISSN 0084-6198.

Lindsey:1971:SRD

[299] C. H. Lindsey. Sliced row-displays and common sub-expressions. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):22–24, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:IWRa

[300] Peter Naur. IFIP WG2.1: Report of subcommittee on data processing and transput. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):25–39, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:IWRb

[301] Peter Naur. IFIP WG2.1: Report of subcommittee on maintenance of and improvements to ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):40–47, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:IWRc

[302] Peter Naur. IFIP WG2.1: Report of subcommittee on conversational languages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32): 48–49, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:IWRd

[303] Peter Naur. IFIP WG2.1: Report of subcommittee on interface between languages and operating systems. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):49–50, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1971:ARN

[304] Peter Naur. Activity report no. 10 & 11. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (32):51–56, May 1971. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:EN

[305] Peter Naur. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):1–2, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:A

[306] Peter Naur. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):3–4, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:C

[307] Peter Naur. Correspondence. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):4, March 1972. CODEN ALGOBG. ISSN 0084-6198.

vanWingen:1972:PSI

[308] J. W. van Wingen. A proposal for a simple input–output system in AL-GOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33): 5–16, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Roberts:1972:QRA

[309] J. D. Roberts. On the question of referencing active processes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):17–20, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:IWSa

[310] Peter Naur. IFIP WG2.1 subcommittee: Maintenance of and improvements to ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):21–25, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:IWSb

[311] Peter Naur. IFIP WG2.1 subcommittee: Data processing and transput. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):26–42, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:IWSc

[312] Peter Naur. IFIP WG2.1 subcommittee: Sublanguages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):43–46, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Naur:1972:IWSd

[313] Peter Naur. IFIP WG2.1 subcommittee: Operating systems and conversational languages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):46, March 1972. CODEN ALGOBG. ISSN 0084-6198.

Bourne:1972:SIA

[314] S. R. Bourne and M. J. T. Guy. Suggestions for improvements to ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):47–53, March 1972. CODEN ALGOBG. ISSN 0084-6198. See comments [315].

Bourne:1972:CSI

[315] S. R. Bourne and M. J. T. Guy. Comments on suggested improvements to ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (33):54–58, March 1972. CODEN ALGOBG. ISSN 0084-6198. See [314].

Lindsey:1972:EN

[316] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (34):1–2, July 1972. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1972:RMW

[317] C. H. Lindsey. Report on the meeting of WG2.1 at Fontainebleau. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (34):3–5, July 1972. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1972:RCI

[318] C. H. Lindsey. Report on considered improvements (Fontainebleau 9). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (34):6–27, July 1972. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1972:PRT

[319] C. H. Lindsey. Proposals for revision of the transput sections of the Report (Fontainebleau 10). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (34):28–40, July 1972. CODEN ALGOBG. ISSN 0084-6198.

Freeman:1972:SRC

[320] W. Freeman. Suggestions regarding certain representations in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (34):41–44, July 1972. CODEN ALGOBG. ISSN 0084-6198.

Rain:1972:SFL

[321] Mark Rain. Some formal language aspects of MARY. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (34):45–81, July 1972. CODEN ALGOBG. ISSN 0084-6198.

vanGils:1972:PEL

[322] A. J. M. van Gils. The philosophy of an extensible language. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (34):82–88, July 1972. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:ENa

[323] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):2–3, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:Aa

[324] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):4, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:FRIa

[325] C. H. Lindsey. Further report on improvements to ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):5–13, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Hibbard:1973:MGP

[326] P. G. Hibbard. A minimum general purpose sublanguage of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):14–23, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Fokkinga:1973:SRP

[327] M. M. Fokkinga. A self-reproducing program. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):24–26, March 1973. CODEN AL-GOBG. ISSN 0084-6198.

Bahrs:1973:AAN

[328] A. A. Bahrs, A. P. Ershov, and A. F. Rar. Adaption of ALGOL 68 to national languages and implementation peculiarities. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):27–28, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Wegner:1973:GCA

[329] Eberhard Wegner. A generalized completer for ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):29–31, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Moudry:1973:RRD

[330] J. Moudry, J. Kral, J. Nadrchal, and J. Sokol. Recognition of routinedenotations in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):32–33, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Kral:1973:EME

[331] J. Kral. The equivalence of modes and the equivalence of finite automata. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35): 34–35, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Rain:1973:DGV

[332] Mark Rain. Declaration groups and values of declarations in MARY. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (35):36–37, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Rain:1973:OEM

[333] Mark Rain. Operation expressions in MARY. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):38–44, March 1973. CODEN AL-GOBG. ISSN 0084-6198.

Meertens:1973:SSC

[334] L. Meertens. On static scope checking in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (35):45–58, March 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:ENb

[335] C. H. Lindsey. Editors’ note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):1–3, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:Ab

[336] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):3–6, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:C

[337] C. H. Lindsey. Correspondence. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):7, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1973:FRIb

[338] C. H. Lindsey. Final report on improvements to ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):8–16, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Boom:1973:NBA

[339] H. Boom. Note on balancing in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):17–24, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Levinson:1973:SA

[340] M. R. Levinson. Simulation with ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):25–27, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Sintzoff:1973:RAR

[341] M. Sintzoff. On the revised ALGOL 68 Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (36):28–39, November 1973. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:ENa

[342] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):2, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:Aa

[343] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):3, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:WSA

[344] C. H. Lindsey. WG 2.1 subcommittee on ALGOL 68 support. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):4–11, July 1974. CO-DEN ALGOBG. ISSN 0084-6198.

Schuman:1974:TMP

[345] S. A. Schuman. Toward modular programming in high-level languages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):12–23, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:PP

[346] C. H. Lindsey. Partial parametrization. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):24–25, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:M

[347] C. H. Lindsey. Modals. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):26–29, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Hibbard:1974:PSA

[348] P. G. Hibbard. A proposed sublanguage of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):30–53, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Sintzoff:1974:BRA

[349] M. Sintzoff. A brief review of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):54–62, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Boom:1974:EUA

[350] H. Boom. Experience with the use of ALGOL as SIL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):63–67, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Peck:1974:ASC

[351] J. E. L. Peck, M. Sintzoff, and J. M. Watt. ALGOL 68 syntax chart. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (37):68, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:RARa

[352] C. H. Lindsey. Revised ALGOL 68 Report. ERRATA-2. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (37):69–79, July 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:ENb

[353] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):2, December 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:Ab

[354] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):3–4, December 1974. CODEN ALGOBG. ISSN 0084-6198.

DeMorgan:1974:CAR

[355] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. A commentary on the ALGOL 60 Revised Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):5–38, December 1974. CODEN ALGOBG. ISSN 0084-

Shearn:1974:VSA

[356] D. C. S. Shearn. A view on simulation in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):39–42, December 1974. CODEN ALGOBG. ISSN 0084-6198.

Levinson:1974:SA

[357] M. R. Levinson. Simulation with ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):43–44, December 1974. CODEN ALGOBG. ISSN 0084-6198.

Feldmann:1974:IMR

[358] Harry Feldmann. An interpretation for making references (in ALGOL 68). ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):45–51, December 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:RARb

[359] C. H. Lindsey. Revised ALGOL 68 Report. ERRATA-3. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):52–55, December 1974. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1974:QIP

[360] C. H. Lindsey. Questionnaire to implementors on the proposed revision to ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (38):56–57, December 1974. CODEN ALGOBG. ISSN 0084-6198.

Hibbard:1976:SA

[361] P. G. Hibbard. A sublanguage of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 39):1–15, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:ENa

[362] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):2–3, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:Aa

[363] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):4–5, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:SPP

[364] C. H. Lindsey. Specification of partial parametrization proposal. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):6–9, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Wichmann:1976:SRA

[365] B. A. Wichmann. A summary of the replies to the ALGOL 60 questionnaire. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39): 10–12, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Currie:1976:MPA

[366] I. F. Currie. Modular programming in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):13–19, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:PMF

[367] C. H. Lindsey. Proposal for a modules facility in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):20–29, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Meertens:1976:NID

[368] L. G. L. T. Meertens. A note on integer division. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):30–32, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:RAR

[369] C. H. Lindsey. Revised ALGOL 68 Report. ERRATA-4. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (39):33–38, February 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:ENb

[370] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):2, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:Ab

[371] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):3–5, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1976:LE

[372] C. H. Lindsey. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):6–7, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Hilden:1976:IDO

[373] J. Hilden. Integral division once more. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):8–9, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Sklenar:1976:MII

[374] Ivan Sklenar. A method of implementation of independently compiled routine texts in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):10–20, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Haetjens:1976:PSS

[375] R. Haetjens. Proposal for a simple syntax for the ALGOL 68 unit. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):21–23, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Hansen:1976:IDR

[376] Wilfred J. Hansen and Hendrik Boom. IFIP document — report on the standard hardware representation for ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (40):24–43, August 1976. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1977:EN

[377] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):2, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1977:A

[378] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):3–6, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1977:LE

[379] C. H. Lindsey. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):7, July 1977. CODEN ALGOBG. ISSN 0084-6198.

DeMorgan:1977:A

[380] R. De Morgan. The Algollers. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):8–9, July 1977. CODEN ALGOBG. ISSN 0084-6198.

VanVliet:1977:ATC

[381] J. C. Van Vliet. On the ALGOL 68 transput conversion routines. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):10–24, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Holdsworth:1977:VTP

[382] D. Holdsworth. Visibility and teachability of I/O processing in high level languages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41): 25–39, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Walker:1977:SAP

[383] A. N. Walker. The syntax of an ALGOL program. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):40–46, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Bell:1977:TRS

[384] R. Bell. A token-recognizer for the standard harware representation of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):47–70, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Hunter:1977:SAC

[385] R. B. Hunter, R. Kingslake, and A. D. McGettrick. Some ALGOL 68 compilers. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41): 71–73, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1977:ASE

[386] C. H. Lindsey. ALGOL 60 supplement — errata. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):74, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1977:RRE

[387] C. H. Lindsey. Revised report — errata. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (41):74, July 1977. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:ENa

[388] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):2, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:Aa

[389] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):3–5, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:LEa

[390] C. H. Lindsey. Letter to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):5, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Nadrchal:1978:IT

[391] J. Nadrchal. Implementation on TESLA 200. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):5, May 1978. CODEN AL-

GOBG. ISSN 0084-6198. URL https://archive.computerhistory.
org/resources/text/algol/algol_bulletin/A42/P21.HTM.

Lindsey:1978:CRRa

[392] C. H. Lindsey. Commentaries on the Revised Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):6, May 1978. CODEN AL-GOBG. ISSN 0084-6198.

Lindsey:1978:PET

[393] C. H. Lindsey. Parameterisation of the environment for transportable numerical software. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):7–10, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Hansen:1978:TSS

[394] Wilfred J. Hansen. Trouble spots in the standard hardware representation for ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):11–13, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Hansen:1978:AHR

[395] Wilfred J. Hansen. ALGOL 68 hardware representation recommendations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):14–16, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Bell:1978:CDT

[396] R. Bell. Corrections to and discussion of “A Token-recognizer for the Standard Harware Representation of ALGOL 68”. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):17–21, May 1978. CODEN AL-GOBG. ISSN 0084-6198.

Lindsey:1978:AYF

[397] C. H. Lindsey. ALGOL 68 and your friendly neighbourhood operating system. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42): 22–35, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Black:1978:PAH

[398] A. P. Black and V. J. Rayward-Smith. Proposals for ALGOL H — a superlanguage of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):36–49, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Baker:1978:MCF

[399] J. P. Baker. The most contrived factorial program. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):50–52, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Pemberton:1978:GAA

[400] Steven Pemberton. Grammar analysis with ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):53–55, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Geurts:1978:RA

[401] Leo Geurts and Lambert Meertens. Remarks on abstracto. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):56–63, May 1978. CO-DEN ALGOBG. ISSN 0084-6198.

Dewar:1978:APA

[402] R. Dewar and J. Schwartz. ‘Abstracto’ Project for an algorithm specification language. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):64–73, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Sintzoff:1978:LDP

[403] Michel Sintzoff. On language design for program construction. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):74–84, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Haentjens:1978:ARF

[404] R. Haentjens and P. E. Gennart. ALGOL 68 (revised) format-text syntax chart. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (42):85, May 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:ENb

[405] C. H. Lindsey. Editor’s note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):1–2, December 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:Ab

[406] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):2–3, December 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:LEb

[407] C. H. Lindsey. Letter to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):4–5, December 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:CRRb

[408] C. H. Lindsey. Commentaries on the Revised Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):6–18, December 1978. CO-DEN ALGOBG. ISSN 0084-6198.

Lindsey:1978:MSC

[409] C. H. Lindsey and H. J. Boom. A modules and separate compilation facility for ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):19–53, December 1978. CODEN ALGOBG. ISSN 0084-6198. See comments [436] and errata [442].

King:1978:MBP

[410] P. R. King, G. Cormack, G. Dueck, N. Jung, G. Kusner, and J. Melnyk. MABEL: A beginner’s programming language. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):54–83, December 1978. CODEN ALGOBG. ISSN 0084-6198.

Sintzoff:1978:PFU

[411] M. Sintzoff. Properties, feasibility and usefulness of a language for programming programs. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):84–90, December 1978. CODEN ALGOBG. ISSN 0084-6198.

Boom:1978:AT

[412] Hendrik Boom. ABSTRACTd thoughts. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (43):91–96, December 1978. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1979:EN

[413] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):2, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1979:A

[414] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):2–5, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1979:CRR

[415] C. H. Lindsey. Commentaries on the Revised Report. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):6–7, May 1979. CODEN AL-GOBG. ISSN 0084-6198.

Wupper:1979:EAT

[416] Hanno Wupper. Experiences with ALGOL 68 transput. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):8–14, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Cheney:1979:CLL

[417] C. J. Cheney, C. H. Lindsey, L. Meertens, and H. Wupper. Changing line lengths in random files. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):15, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Jonkers:1979:FSL

[418] H. B. M. Jonkers. A finite state lexical analyzer for the standard hardware representation of ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):16–51, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Rayward-Smith:1979:UAP

[419] V. J. Rayward-Smith. The use of ALGOL 68 pattern matching to describe a formal logic system. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):52–62, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Ajueva:1979:SRD

[420] Vera Ajueva, A. N. Maslov, and V. B. Yakovlev. A schema for reading data in formatless input. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (44):63, May 1979. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:ENa

[421] C. H. Lindsey. Editor’s note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):1–2, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:Aa

[422] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):2–3, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Reid:1980:FMF

[423] J. K. Reid. Functions for manipulating floating-point numbers. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):6–15, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:AIF

[424] C. H. Lindsey. ALGOL 68 implementations — FLACC. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):16, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:AIA

[425] C. H. Lindsey. ALGOL 68 implementations — ALGOL 68C — release 1. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):17–22, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Baszenski:1980:PCT

[426] G. Baszenski and H. Wupper. A proposal for conversational transput. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):23–29, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Grune:1980:OA

[427] D. Grune and C. H. Lindsey. Overprinting in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):30–32, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Ru-qian:1980:TAC

[428] Lu Ru-qian. The translation of ALGOL 68 into Chinese. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):33–37, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Ince:1980:AAM

[429] D. C. Ince. ALGOL 68 and algebraic manipulation. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):38–44, January 1980. CO-DEN ALGOBG. ISSN 0084-6198.

Ince:1980:AEL

[430] D. C. Ince. An algorithm for the execution of limited entry decision tables in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (45):45–46, January 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:ENb

[431] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (46):2, December 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:Ab

[432] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (46):2–5, December 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:LE

[433] C. H. Lindsey. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (46):5–6, December 1980. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1980:AII

[434] C. H. Lindsey. ALGOL 68 implementations — the ICL 2900 compiler. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (46):7–8, December 1980. CODEN ALGOBG. ISSN 0084-6198.

Ehlich:1980:MRA

[435] H. Ehlich and H. Wupper. More remarks on ALGOL 68 transput. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (46):9–11, December 1980. CODEN ALGOBG. ISSN 0084-6198.

Cowling:1980:CPM

[436] A. J. Cowling. Comments on the proposals for modules and separate compilation facilities in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (46):12–26, December 1980. CODEN ALGOBG. ISSN 0084-6198. See [409].

Lindsey:1980:AI

[437] C. H. Lindsey. An ALGOL 68 indenter. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (46):27–35, December 1980. CODEN ALGOBG. ISSN 0084-6198.

vanWijngaarden:1981:RRA

[438] A. van Wijngaarden. Revised report of the algorithmic language Algol 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (Sup 47):1–119, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1981:EN

[439] C. H. Lindsey. Editor’s note. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):2, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1981:A

[440] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):2–7, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1981:LE

[441] C. H. Lindsey. Letters to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):7–10, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1981:EMS

[442] C. H. Lindsey. Errata to ‘A modules and Separate Compilation Facility for ALGOL 68’. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):11–12, August 1981. CODEN ALGOBG. ISSN 0084-6198. See [409].

Lindsey:1981:EAT

[443] C. H. Lindsey. Errata to ‘ALGOL 68 transput, pt II: An implementation model’. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47): 12–14, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1981:SVA

[444] C. H. Lindsey. Survey of viable ALGOL 68 implementations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):15–18, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Thomson:1981:SRP

[445] C. M. Thomson. Self-replicating programs and n-cycle programs. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):19–20, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Thomson:1981:ALL

[446] C. M. Thomson. ALGOL 68 as a living language. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):21–24, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Levinson:1981:AIC

[447] M. R. Levinson. An ALGOL 68 implementation companion. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (47):25–26, August 1981. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1982:EN

[448] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (48):2, August 1982. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1982:A

[449] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (48):2–6, August 1982. CODEN ALGOBG. ISSN 0084-6198.

Meertens:1982:QRB

[450] L. G. L. T. Meertens. Quick reference to B. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (48):7–17, August 1982. CODEN ALGOBG. ISSN 0084-6198.

Currie:1982:MCS

[451] I. F. Currie and N. E. Peeling. Modular compilation systems for high level programming languages. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (48):18–26, August 1982. CODEN ALGOBG. ISSN 0084-6198.

Hodgson:1982:NAL

[452] G. S. Hodgson. The NAG ALGOL 68 library. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (48):27–38, August 1982. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1983:ENa

[453] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (49):1–2, May 1983. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1983:Aa

[454] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (49):2–4, May 1983. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1983:LE

[455] C. H. Lindsey. Letter to the Editor. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (49):5–6, May 1983. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1983:CMA

[456] C. H. Lindsey. Clarification to modified ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (49):7–9, May 1983. CODEN AL-GOBG. ISSN 0084-6198.

Lindsey:1983:PEH

[457] C. H. Lindsey. A proposal for exception handling in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (49):10–15, May 1983. CODEN ALGOBG. ISSN 0084-6198.

Thomas:1983:EHM

[458] Martyn Thomas. An exception-handling mechanism for ALGOL 68. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (49):16–17, May 1983. CODEN ALGOBG. ISSN 0084-6198.

Elsworth:1983:SRP

[459] E. F. Elsworth. A self-replicating program in ALGOL 68C. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (49):18–19, May 1983. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1983:ENb

[460] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (50):2, December 1983. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1983:Ab

[461] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (50):3–5, December 1983. CODEN ALGOBG. ISSN 0084-6198.

Owen:1983:ASP

[462] Kenneth Owen. The art and science of programming. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (50):6–9, December 1983. CO-DEN ALGOBG. ISSN 0084-6198.

Allison:1983:EPS

[463] Lloyd Allison. An executable Prolog semantics. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (50):10–18, December 1983. CODEN ALGOBG. ISSN 0084-6198.

Outteridge:1983:LMC

[464] David Outteridge. A library mechanism for the CDC ALGOL 68 compiler. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (50):19–27, December 1983. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1984:EN

[465] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (51):2, December 1984. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1984:A

[466] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (51):3–4, December 1984. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1984:SVA

[467] C. H. Lindsey. Survey of viable ALGOL 68 implementations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (51):5–8, December 1984. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1984:IRA

[468] C. H. Lindsey and I. Domville. Implementing random in ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (51):9–15, December 1984. CODEN ALGOBG. ISSN 0084-6198.

Craven:1984:IA

[469] Peter G. Craven. Interactive ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (51):16–18, December 1984. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1988:EN

[470] C. H. Lindsey. Editor’s notes. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (52):1–2, August 1988. CODEN ALGOBG. ISSN 0084-6198.

Anonymous:1988:LAB

[471] Anonymous. The last Algol Bulletin. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (52):1–38, August 1988. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1988:A

[472] C. H. Lindsey. Announcements. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (52):2–4, August 1988. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1988:SVA

[473] C. H. Lindsey. Survey of viable ALGOL 68 implementations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (52):5–8, August 1988. CODEN ALGOBG. ISSN 0084-6198.

Hackenberg:1988:IAC

[474] Klaus Hackenberg. Implementation of ALGOL 68 on the Cyber 205. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), (52):9–13, August 1988. CODEN ALGOBG. ISSN 0084-6198.

Tseytin:1988:EHP

[475] G. S. Tseytin. An exception handling proposal for ALGOL 68. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (52):14–26, August 1988. CODEN ALGOBG. ISSN 0084-6198.

Lindsey:1988:BTS

[476] C. H. Lindsey. A browse through some early Bulletins. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (52):27–38, August 1988. CODEN ALGOBG. ISSN 0084-6198.